# Reward Inside the Model: A Lightweight Hidden-State Reward Model for LLM’s Best-of-N sampling

### v1

@@ C.3 ELHSR Training Details
For each dataset, we began by sampling 6500 unique instances. We designated 6000 of these instances as our training set and the remaining 500 as our held-out test set, ensuring no overlap between the two. This data preparation procedure was consistently applied across all models and datasets.
We employed early stopping [Yao et al., 2007] to prevent overfitting. We further split the training data into training and validation sets using an 80/20 ratio. After each training epoch, we evaluated the model’s performance on the validation set. We used the cross-entropy with logits loss on validation set as the criterion for early stopping. We used a patience of 3 epochs, meaning that training was terminated if the validation loss did not improve for 3 consecutive epochs. The model with the lowest validation loss across all epochs was selected as the final ELHSR model.
C.4 Generation Details
For generating reasoning paths, we employed bfloat16 precision [Burgess et al., 2019] to accelerate inference and reduce memory consumption, using vLLM library [Kwon et al., 2023] and Huggingface’s transformers library [Wolf et al., 2019]. For each of the 6000 training instances, we generated 8 reasoning paths (rollouts), with max_new_tokens to be 1024. We set temperature to be 1.0, but set top_p to be 0.9 to prevent generating garbage text. The input format strictly followed the instruction-question template:
@@ C.4 Generation Details
C.5 Evaluation Details
Due to the potential for equivalent but syntactically different answers (e.g., "6/5" and "1.2"), a simple string comparison is insufficient for determining correctness in the MATH and GSM8K datasets. Therefore, we adopted a more sophisticated evaluation approach, leveraging the implementation from the code repository of Yuan et al. [2025]. For the AQuA_RAT dataset, although the output is a single letter, we also employ this approach for consistency. This implementation allows for the robust comparison of mathematical expressions, treating semantically equivalent answers as correct. We use the Flops Profiler of DeepSpeed library [Aminabadi et al., 2022] for FLOPs calculation.
C.6 Details on EurusRM-7B fine-tuning
As a baseline to validate the superior performance of our ELHSR method, we fine-tuned the EurusRM-7B reward model [Yuan et al., 2025] to provide a more comprehensive comparison, although it has already been pretrained on MATH and GSM8K dataset, which may give it a potential advantage. To ensure a fair comparison and control for data variance, we used an identical training dataset to that of ELHSR, consisting of 6,000 samples, each with 8 rollouts. Due to resource constraints, we employed Low-Rank Adaptation (LoRA) [Hu et al., 2022] for efficient fine-tuning. The AdamW optimizer was used with the following hyperparameters:
  • Learning Rate: $1\times 10^{-5}$
  • Weight Decay: $1\times 10^{-5}$
  • Batch Size: 1
Appendix D Results on the CRUXEval-O Dataset
In addition to the MATH, GSM8K, and AQuA_RAT datasets, we also evaluated ELHSR’s performance on the CRUXEval-O dataset [Gu et al., 2024a] to assess its capabilities in code comprehension. Our experimental setup is detailed below.
For training ELHSR and fine-tuning EurusRM-7B, we utilized 600 instances from CRUXEval-O. A separate set of 200 instances was used for evaluation, applied to both ELHSR and all baseline models. The instruction template employed for the CRUXEval-O dataset was:
@@ Appendix D Results on the CRUXEval-O Dataset
[…]
@@
Following answer extraction, we employed Python’s ast.literal_eval function to parse both the model’s output and the reference answer, and subsequently compared the parsed values. In cases where parsing failed, we implemented a fallback mechanism: treating the extracted content as a string, removing any surrounding quotes, and then comparing it with the reference answer (also stripped of surrounding quotes).
Table 5 presents the results of our evaluation. The data demonstrates that ELHSR outperforms the baselines on the code comprehension task as well, even trained only on a very small training set (600 instances). This further supports the effectiveness of our approach.
Table 5: Different methods’ best-of-N sampling performance on CRUXEval-O dataset. The training data size is 600 samples. Results represent accuracy (percentages omitted). Our method outperforms baselines. Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
CRUXEval-O Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 15.0 BoN@1: 36.5 BoN@1: 51.5 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 12.5 20.0 19.0 44.0 41.5 40.0 54.5 56.5 57.0 38.3 Skywork-Llama3.1-8B 27.0 39.0 42.5 49.5 52.5 48.0 55.5 56.5 57.5 47.6 Starling-7B 11.5 11.5 14.5 39.5 32.5 38.0 53.0 54.0 50.0 33.8 Ultra-13B 21.5 29.5 36.5 49.0 51.5 53.5 55.5 53.0 56.0 45.1 Fine-tuned Eurus-7B 25.5 38.0 37.5 50.5 56.0 55.5 58.0 59.5 60.0 48.9 ELHSR (ours) 29.0 37.0 38.5 58.5 57.0 57.0 57.0 61.5 58.0 50.4
@@ Appendix E Ablation Study
E.1 Impact of Loss Function
This section investigates the impact of different loss functions on the performance of the ELHSR model. We compare our chosen cross-entropy with logits loss with hinge loss, Direct Preference Optimization (DPO) loss, InfoNCA loss, and NCA loss.
E.1.1 Loss Function Formulations
  • Cross-Entropy with Logits Loss [Rafailov et al., 2024]: Given a binary label $y\in\{0,1\}$ (where 1 represents a correct path and 0 an incorrect path) and a predicted reward $R$, the CE loss with logits is:
    $\mathcal{L}_{CE}=-[y\log(\sigma(R))+(1-y)\log(1-\sigma(R))]$ (12)
@@ E.1.1 Loss Function Formulations
  • Hinge Loss [Rosasco et al., 2004]: As defined in Section 3.2, the hinge loss for a reasoning path with a predicted reward $R$ is:
    $\mathcal{L}_{hinge}=\max\left(0,1-y\cdot R\right)$ (13)
  • Direct Preference Optimization (DPO) Loss [Rafailov et al., 2023]: Given the reward for a positive (chosen) path $r^{+}$ and the reward for a negative (rejected) path $r^{-}$, the DPO loss is:
    $\mathcal{L}_{DPO}=-\log(\sigma(r^{+}-r^{-}))$ (14)
@@
  • InfoNCA Loss [Chen et al., 2024]: Given predicted rewards $r_{pred}\in\mathbb{R}^{K}$ for $K$ candidate paths and corresponding ground-truth rewards $r_{gt}\in\mathbb{R}^{K}$, and a temperature parameter $\alpha$, the InfoNCA loss is:
    $\mathcal{L}_{InfoNCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log\left(\text{softmax}(r_{pred})_{i}\right)\right]$ (15)
  • NCA Loss [Chen et al., 2024]: With the same notation as InfoNCA, the NCA loss is:
    $\mathcal{L}_{NCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log(\sigma(r_{pred_{i}}))+\frac{1}{K}\log(\sigma(-r_{pred_{i}}))\right]$ (16)
E.1.2 Experimental Results
We train the ELHSR model using each of these loss functions, keeping all other hyperparameters constant (as detailed in Appendix A). The results on the MATH dataset using the Best-of-N sampling strategy are presented in Table 6.
Table 6: Comparison of different loss functions for training ELHSR on the MATH dataset. Bold: best performance; Underline: second-best performance.
@@ E.1.2 Experimental Results
E.2 Ablation Study on the Gating Mechanism
To assess the necessity of the gating mechanism within our ELHSR framework, we conducted an ablation study on the MATH dataset. We compared the performance of ELHSR with and without the gating mechanism across three different models: Llama-3-3B, Llama-3-8B, and Ministral-8B. The results are summarized in Table 7.
Table 7: Ablation Study on the Gating Mechanism on the MATH Dataset. Bold: best performance;
@@ E.2 Ablation Study on the Gating Mechanism
The results indicate that while ELHSR without the gating mechanism can achieve reasonable performance, it consistently underperforms ELHSR with the gating mechanism across all models and k values. This suggests that the gating mechanism plays a crucial role in improving reasoning accuracy.
Appendix F Examples of ELHSR Scored Response
@@
We provide specific examples of ELHSR scores for LLM reasoning paths in Table 8. We highlight the top and bottom 10 tokens based on their ELHSR gating $\times$ reward values, while also preserving special tokens such as <|eot_id|> (end-of-turn identifier). The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen (approximating green), while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon (approximating red). Each highlighted token is also annotated with its corresponding gating $\times$ reward value. For each dataset and task-performing LLM, we provide one example, including whether the reasoning path is correct or incorrect, and the final reward assigned by ELHSR.
@@
Table 8 presents nine examples of ELHSR’s scored reasoning paths. Notably, ELHSR’s gating $\times$ reward scores demonstrate sensitivity to specific elements within the reasoning sequence, tending to assign high scores (highlighted in limegreen) to numerical and mathematical symbols, as well as tokens like “answer”, “boxed”, and the special token ‘<|eot_id|>’ when the overall reward is positive, and low scores (highlighted in salmon) to these same kinds of tokens when the overall reward is negative. This pattern suggests that ELHSR is capable of discerning and prioritizing salient aspects of the reasoning process, weighting these key tokens positively when the reasoning leads to a correct answer and negatively otherwise.
@@
Table 8: Examples of ELHSR Scored Reasoning Paths. The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen, while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon. Each highlighted token is also annotated with its corresponding gating $\times$ reward value. Special tokens such as <|eot_id|> (end-of-turn identifier) are preserved.
@@
Example 4 Dataset GSM8K Task-performing LLM Llama-3.2-3B Question A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take? Response To solve the problem, we first need to determine the amount of blue and white fiber required. Since it takes 2 bolts of blue fiber, we can start by writing that as 2. Let’s denote the amount of white fiber as x. According to the problem, it takes half that much white fiber, so the equation for the total amount of fiber required would be: 2 (blue) + -0.150.5x (white) = total Since the problem doesn’t specify the total amount of fiber, let’s assume that we are looking for the total amount of fiber in terms of x-0.37. The total amount of fiber required is the sum of blue and white0.02 fibers, which is 2 + 0.5x. Since we can’t simplify this expression further, the solution is 0.012 + 0-0.11.5x-0.27. However, to adhere to the requested answer format, the final answer should be presented in the box and include the value of x. In this case, since x is an unknown value and not a numerical one, it cannot be simplified to a single0.00 number, the solution is expressed in terms of x,0.00 2 + -0.180.5x,0.00 but it is not exactly fitting the requested answer format, that being a simple0.02 answer in a box. Since it does not specify the value of x, it cannot be simplified further in terms of a single number. But since it is required to follow the format to the letter the answer is: \0.01boxed4.23{2.542-8.06 + 0-11.09.5-0.36x-12.63}0.00<|eot_id|>-5.31 Final Reward -5.54 Correctness False Reference Answer $3$
@@ Appendix G Additional Results
This section provides additional results of our experiments.
Table 9: Training ELHSR solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency.
@@
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
@@
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 76.0 79.6 81.0 89.2 90.6 91.2 87.4 87.0 90.2 85.8
@@
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 58.8 63.4 65.2 68.6 77.2 78.0 69.6 74.4 74.0 69.9
Figure 6: Comparison of average time and FLOPs per sample for the proposed ELHSR reward model and baseline reward models, evaluated on different datasets and task-performing models. The y-axis is plotted in log-scale. The results show that ELHSR achieves orders-of-magnitude higher efficiency than the baselines.
Figure 7: ELHSR has positive scaling with increased number of training samples and with the number of reasoning paths for inference.
Figure 8: Performance of combined ELHSR and external reward models across different datasets and task-performing models, using both rank selection and scaled averaging.
Appendix H Broader Impacts and Safeguards
The development of ELHSR presents both opportunities and challenges for society. On the positive side, its efficiency could democratize access to high-performing LLMs, enabling smaller labs, startups, and academic groups to iterate more rapidly on large language model (LLM) applications without requiring massive compute budgets. In addition, its remarkable efficiency can also promote more sustainable AI practices through reduced computational demands, cutting energy consumption and carbon footprint across the machine-learning lifecycle. However, the efficiency of ELHSR might also lower the bar for malicious use, possibly towards the development of harmful applications like disinformation attacks.
To mitigate these harms, we recommend proactive measures in the form of robust security measures, transparent reporting and auditing, clearly defined ethical guidelines for responsible use, and increased investments in AI safety and alignment research. By critically analyzing and taking into account these broader social implications, we can strive to ensure that ELHSR and other similar technologies are deployed to bring maximum benefit with minimal potential harm.
Appendix I Limitations
Admittedly, this work has some limitations. ELHSR’s reliance on access to the LLM’s hidden states or logits makes it potentially less applicable to commercial LLMs that do not expose these information to users. However, the efficiency and ease of deployment with ELHSR ensure that it is well-suited to LLM providers themselves. Meanwhile, although the preliminary experiment in Section 3.1 motivates the linearity of LLM hidden states and Appendix F provides analysis of ELHSR on some examples, future work could benefit from exploring deeper theoretical analysis of ELHSR.
Appendix J Licenses and Copyrights Across Assets
This section provides a comprehensive disclosure of the citations, links, and licenses pertaining to the assets utilized in this research. We affirm our adherence to the stipulated license terms. We extend our sincere gratitude to the respective asset owners for their contributions.
J.1 Models
  1. 1.
    Llama-3.2-3B-Instruct
  2. 2.
    Llama-3.1-8B-Instruct
  3. 3.
    Ministral-8B-Instruct
  4. 4.
    EurusRM-7B
  5. 5.
    Skywork-Reward-Llama-3.1-8B-v0.2
  6. 6.
    Starling-7B
  7. 7.
    UltraRM-13B
J.2 Datasets
  1. 1.
    MATH
  2. 2.
    GSM8K
  3. 3.
    AQuA_RAT
  4. 4.
    CRUXEval
J.3 Libraries
  1. 1.
    PyTorch
  2. 2.
    scikit-learn
  3. 3.
    vLLM
  4. 4.
    Transformers (Hugging Face)
  5. 5.
    DeepSpeed
[…]
[appendices]

### v3

@@ D.3. SWIFT Training Details
For each dataset, we created a training set of 6000 instances, drawn from the original training set, and a held-out test set of 500 instances, drawn from the original validation set (due to the absence of labels in some datasets’ test sets). This consistent data preparation approach was applied across all datasets, with the exceptions of GSM8K (where the original test set was used due to the lack of a validation set) and Imbue Code Comprehension (where we performed a manual train/test split as no split was provided).
In addition, we need to generate reasoning path and determine whether each reasoning path is correct. The generation details are provided in Appendix D.4 and the method of determining each reasoning path’s correctness is detailed in Appendix D.5.
We employed early stopping (Yao et al., 2007) to prevent overfitting. We further split the training data into training and validation sets using an 80/20 ratio. After each training epoch, we evaluated the model’s performance on the validation set. We used the cross-entropy with logits loss on validation set as the criterion for early stopping. We used a patience of 3 epochs, meaning that training was terminated if the validation loss did not improve for 3 consecutive epochs. The model with the lowest validation loss across all epochs was selected as the final SWIFT model.
D.4. Generation Details
For generating reasoning paths, we employed bfloat16 precision (Burgess et al., 2019) to accelerate inference and reduce memory consumption, using vLLM library (Kwon et al., 2023) and Huggingface’s transformers library (Wolf et al., 2019). For each of the 6000 training instances, we generated 8 reasoning paths (rollouts), with max_new_tokens to be 1024. We set temperature to be 1.0, but set top_p to be 0.9 to prevent generating garbage text. The input format strictly followed the instruction-question template:
@@ D.4. Generation Details
  • Imbue Code Comprehension:
    Analyze the following problem step-by-step. The question includes a list of choices.
@@
[…]
  • HellaSwag:
    You are given a context and four possible continuations (A--D). Decide which option best continues the context.
[…]
  • CoinFlip:
    Answer the following Boolean question with detailed reasoning.
[…]
D.5. Evaluation Details / Method to identify correctness
This section details how we determine each reasoning path’s correctness for both training and evaluation, and the method for FLOPs calculation.
Due to the potential for equivalent but syntactically different answers (e.g., "6/5" and "1.2"), a simple string comparison is insufficient for determining correctness in the MATH and GSM8K datasets. Therefore, we adopted a more sophisticated evaluation approach, leveraging the implementation from the code repository of Yuan et al. (2025). For the AQuA_RAT, HellaSwag and CoinFlip dataset, although the output is a single letter, we also employ this approach for consistency. This implementation allows for the robust comparison of mathematical expressions, treating semantically equivalent answers as correct.
@@ D.5. Evaluation Details / Method to identify correctness
  • Removed any occurrences of $\text{…}$‘ by replacing them with the inner text.
  • Replaced escaped braces $\{$‘ and ’\}‘ with regular braces ’{‘ and ’}‘.
@@
Following answer extraction, we employed Python’s
ast.literal_eval function to parse both the model’s output and the reference answer, and subsequently compared the parsed values. In cases where parsing failed, we implemented a fallback mechanism: treating the extracted content as a string, removing any surrounding quotes, and then comparing it with the reference answer (also stripped of surrounding quotes).
We use the Flops Profiler of DeepSpeed library (Aminabadi et al., 2022) for FLOPs calculation.
D.6. Details on EurusRM-7B fine-tuning
As a baseline to validate the superior performance of our SWIFT method, we fine-tuned the EurusRM-7B reward model (Yuan et al., 2025) to provide a more comprehensive comparison, although it has already been pretrained on MATH and GSM8K dataset, which may give it a potential advantage. To ensure a fair comparison and control for data variance, we used an identical training dataset to that of SWIFT, consisting of 6,000 samples, each with 8 rollouts. Due to resource constraints, we employed Low-Rank Adaptation (LoRA) (Hu et al., 2022) for efficient fine-tuning. The AdamW optimizer was used with the following hyperparameters:
  • Learning Rate: $1\times 10^{-5}$
  • Weight Decay: $1\times 10^{-5}$
  • Batch Size: 1
Our fine-tuning process employs a loss function identical to that of the original paper Yuan et al. (2025). The specific loss function, $\mathcal{L}_{\text{ULTRAINTERACT}}$, is defined as follows:
[…]
In this formulation, $x$ denotes the problem prompt, $y_{c}$ represents the chosen (correct) answer, and $y_{r}$ is the rejected (incorrect) answer. The reward model, parameterized by $\theta$, is denoted by $r_{\theta}(\cdot)$, and $\sigma$ is the sigmoid function. The $\mathcal{L}_{\text{BT}}$ term is a Bradley-Terry-style loss that optimizes the relative rewards, encouraging the reward of the correct answer to be higher than that of the incorrect one. The $\mathcal{L}_{\text{DR}}$ term directly regularizes the rewards by increasing the reward for the chosen answer and decreasing it for the rejected answer.
Furthermore, our data sampling methodology is analogous to the approach described in the original paper. For each problem instance, a training pair is constructed by randomly matching one correct solution with one incorrect solution from the same problem.
@@ Appendix E Ablation Study
E.1. Impact of Loss Function
This section investigates the impact of different loss functions on the performance of the SWIFT model. We compare our chosen cross-entropy with logits loss with hinge loss, Direct Preference Optimization (DPO) loss, InfoNCA loss, and NCA loss.
E.1.1. Loss Function Formulations
  • Cross-Entropy with Logits Loss (Rafailov et al., 2024): Given a binary label $y\in\{0,1\}$ (where 1 represents a correct path and 0 an incorrect path) and a predicted reward $R$, the CE loss with logits is:
    (13) $\mathcal{L}_{CE}=-[y\log(\sigma(R))+(1-y)\log(1-\sigma(R))]$
@@ E.1.1. Loss Function Formulations
  • Hinge Loss (Rosasco et al., 2004): As defined in Section 4.2, the hinge loss for a reasoning path with a predicted reward $R$ is:
    (14) $\mathcal{L}_{hinge}=\max\left(0,1-y\cdot R\right)$
  • Direct Preference Optimization (DPO) Loss (Rafailov et al., 2023): Given the reward for a positive (chosen) path $r^{+}$ and the reward for a negative (rejected) path $r^{-}$, the DPO loss is:
    (15) $\mathcal{L}_{DPO}=-\log(\sigma(r^{+}-r^{-}))$
@@
  • InfoNCA Loss (Chen et al., 2024): Given predicted rewards $r_{pred}\in\mathbb{R}^{K}$ for $K$ candidate paths and corresponding ground-truth rewards $r_{gt}\in\mathbb{R}^{K}$, and a temperature parameter $\alpha$, the InfoNCA loss is:
    (16) $\mathcal{L}_{InfoNCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log\left(\text{softmax}(r_{pred})_{i}\right)\right]$
  • NCA Loss (Chen et al., 2024): With the same notation as InfoNCA, the NCA loss is:
    (17) $\mathcal{L}_{NCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log(\sigma(r_{pred_{i}}))+\frac{1}{K}\log(\sigma(-r_{pred_{i}}))\right]$
E.1.2. Experimental Results
We train the SWIFT model using each of these loss functions, keeping all other hyperparameters constant (as detailed in Appendix A). The results on the MATH dataset using the Best-of-N sampling strategy are presented in Table 7.
Table 7. Comparison of different loss functions for training SWIFT on the MATH dataset. Bold: best performance; Underline: second-best performance.
@@ E.1.2. Experimental Results
E.2. Ablation Study on the Gating Mechanism
To assess the necessity of the gating mechanism within our SWIFT framework, we conducted an ablation study on the MATH dataset. We compared the performance of SWIFT with and without the gating mechanism across three different models: Llama-3-3B, Llama-3-8B, and Ministral-8B. The results are summarized in Table 8.
Table 8. Ablation Study on the Gating Mechanism on the MATH Dataset. Bold: best performance;
@@ E.2. Ablation Study on the Gating Mechanism
The results indicate that while SWIFT without the gating mechanism can achieve reasonable performance, it consistently underperforms SWIFT with the gating mechanism across all models and k values. This suggests that the gating mechanism plays a crucial role in improving reasoning accuracy.
E.3. MATH Sub-domain Breakdown
To better understand when SWIFT helps, we compare SWIFT against Math-Shepherd-7B on MATH and break down the BoN@64 gains by sub-domain (Ministral-8B as the task LLM). Table 9 shows improvements across all sub-domains, with larger gains on categories that typically require more multi-step reasoning.
Table 9. SWIFT gains over Math-Shepherd-7B by MATH sub-domain (Ministral-8B, BoN@64; accuracy difference in percentage points).
Category Gain
Precalculus 7.14
Algebra 8.87
Intermediate Algebra 10.31
Number Theory 9.68
Prealgebra 4.88
Geometry 2.44
Counting & Probability 10.52
Per-category accuracy gains of SWIFT over Math-Shepherd-7B on MATH.
E.4. Layer-wise Ablation Study
Table 10 presents a detailed breakdown of SWIFT’s performance when trained using different subsets of hidden layers from the task-performing LLM. We compare three configurations: using only the first 4 layers, using only the last 4 layers, and using all layers (default). Results are reported for three task LLMs (Llama3-3B, Llama3-8B, Ministral-8B) under BoN@64 on MATH. The first 4 layers yield substantially lower accuracy (38.2/44.6/45.4), confirming that early-layer representations lack sufficient semantic abstraction for reasoning evaluation. In contrast, the last 4 layers achieve 55.4/60.4/61.8, similar to the full model’s performance (53.6/62.6/62.8). This demonstrates that discriminative information for reasoning quality is concentrated in the final layers, allowing SWIFT to remain highly effective even with a reduced layer subset, a useful insight for further efficiency optimization.
Table 10. Layer-wise ablation on MATH (BoN@64). Accuracy (%) across three task LLMs.
Layer Subset Llama3-3B Llama3-8B Ministral-8B
First 4 layers 38.2 44.6 45.4
Last 4 layers 55.4 60.4 61.8
All layers (default) 53.6 62.6 62.8
Appendix F Examples of SWIFT Scored Response
@@
We provide specific examples of SWIFT scores for LLM reasoning paths in Table 11. We highlight the top and bottom 10 tokens based on their SWIFT gating $\times$ reward values, while also preserving special tokens such as <|eot_id|> (end-of-turn identifier). The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen (approximating green), while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon (approximating red). Each highlighted token is also annotated with its corresponding gating $\times$ reward value. For each dataset and task-performing LLM, we provide one example, including whether the reasoning path is correct or incorrect, and the final reward assigned by SWIFT.
@@
Table 11 presents nine examples of SWIFT’s scored reasoning paths. Notably, SWIFT’s gating $\times$ reward scores demonstrate sensitivity to specific elements within the reasoning sequence, tending to assign high scores (highlighted in limegreen) to numerical and mathematical symbols, as well as tokens like “answer”, “boxed”, and the special token ‘<|eot_id|>’ when the overall reward is positive, and low scores (highlighted in salmon) to these same kinds of tokens when the overall reward is negative. This pattern suggests that SWIFT is capable of discerning and prioritizing salient aspects of the reasoning process, weighting these key tokens positively when the reasoning leads to a correct answer and negatively otherwise.
@@
Table 11. Examples of SWIFT Scored Reasoning Paths. The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen, while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon. Each highlighted token is also annotated with its corresponding gating $\times$ reward value. Special tokens such as <|eot_id|> (end-of-turn identifier) are preserved.
@@
Example 4 Dataset GSM8K Task-performing LLM Llama-3.2-3B Question A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take? Response To solve the problem, we first need to determine the amount of blue and white fiber required. Since it takes 2 bolts of blue fiber, we can start by writing that as 2. Let’s denote the amount of white fiber as x. According to the problem, it takes half that much white fiber, so the equation for the total amount of fiber required would be: 2 (blue) + -0.150.5x (white) = total Since the problem doesn’t specify the total amount of fiber, let’s assume that we are looking for the total amount of fiber in terms of x-0.37. The total amount of fiber required is the sum of blue and white0.02 fibers, which is 2 + 0.5x. Since we can’t simplify this expression further, the solution is 0.012 + 0-0.11.5x-0.27. However, to adhere to the requested answer format, the final answer should be presented in the box and include the value of x. In this case, since x is an unknown value and not a numerical one, it cannot be simplified to a single0.00 number, the solution is expressed in terms of x,0.00 2 + -0.180.5x,0.00 but it is not exactly fitting the requested answer format, that being a simple0.02 answer in a box. Since it does not specify the value of x, it cannot be simplified further in terms of a single number. But since it is required to follow the format to the letter the answer is: \0.01boxed4.23{2.542-8.06 + 0-11.09.5-0.36x-12.63}0.00<|eot_id|>-5.31 Final Reward -5.54 Correctness False Reference Answer 33
@@ Appendix G Additional Results
This section provides additional results of our experiments. Please refer to the next page.
Table 12. Training SWIFT solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency.
@@
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
@@
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 76.0 79.6 81.0 89.2 90.6 91.2 87.4 87.0 90.2 85.8
@@
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 58.8 63.4 65.2 68.6 77.2 78.0 69.6 74.4 74.0 69.9
Figure 6. Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on MATH, GSM8K and AQuA_RAT datasets and different task-performing models. The y-axis is plotted in log-scale. The results show that SWIFT achieves orders-of-magnitude higher efficiency than the baselines.
Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on different datasets and task-performing models.
Figure 7. Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on Imbue Code Comprehension, HellaSwag and CoinFlip datasets and different task-performing models. The y-axis is plotted in log-scale. The results show that SWIFT achieves orders-of-magnitude higher efficiency than the baselines.
Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on different datasets and task-performing models.
Figure 8. Performance of combined SWIFT and external reward models across MATH, GSM8K and AQuA_RAT datasets and task-performing models, using both rank selection and scaled averaging.
Performance of combined SWIFT and external reward models across MATH, GSM8K and AQuA_RAT datasets and task-performing models, using both rank selection and scaled averaging.
Figure 9. Performance of combined SWIFT and external reward models across Imbue Code Comprehension, HellaSwag and CoinFlip datasets and task-performing models, using both rank selection and scaled averaging.
Performance of combined SWIFT and external reward models across Imbue Code Comprehension, HellaSwag and CoinFlip datasets and task-performing models, using both rank selection and scaled averaging.